OpenStudy (anonymous):

Would 0=0 be no solution to a system?

OpenStudy (saifoo.khan):

Yes!

OpenStudy (anonymous):

ok, just making sure

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

actually it would be an identity

OpenStudy (anonymous):

It could mean there are an infinite number of solutions to a system

OpenStudy (anonymous):

so it would be all real numbers?

OpenStudy (anonymous):

no its just an identity

OpenStudy (anonymous):

sometimes, a system's solutions space would be a line.

OpenStudy (anonymous):

should i put no solution or all real numbers if my teacher hasnt taught us about identity?

OpenStudy (anonymous):

sometimes it is a plane, sometimes it is a hyperplane.... etc.

OpenStudy (anonymous):

what is the particular system?

OpenStudy (anonymous):

2x+8y=6 -5x-20y=-15

OpenStudy (anonymous):

for that particular system, you get 0 = 0 when you solve it using substitution That means that the two equations in the system represent the same line.

OpenStudy (anonymous):

ok, so do i just put all real numbers?

OpenStudy (anonymous):

Or their slopes are the same. Say "no solution."

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

nope, you put in the equation of the line in slope-intercept form \[SS: y = \frac{1}{4}x + \frac{3}{4}\]

OpenStudy (anonymous):

or am I mistaken?

OpenStudy (amistre64):

0=0 is a true statement, which means that it has solutions

OpenStudy (amistre64):

it just happens to be any and all soultions

OpenStudy (anonymous):

Ah. Right. That's correct. They're the same line.

OpenStudy (anonymous):

yeah I am mistaken\[SS: y = -\frac{1}{4}x + \frac{3}{4}\]

OpenStudy (anonymous):

thanks